What is the next number in the series 7, 2, 0, 7, 8, 4, 3 ?
Please give a step by step solution.

Found 2 solutions by richwmiller, ikleyn:
Answer by richwmiller(17219)

In series we have 7 digits.
7, 2, 0, 7, 8, 4, 3,
1st digit = 7 + 1 = 8 = 5 digit,
2nd digit = 2 + 2 = 4 = 6 digit,
3rd digit = 0 + 3 = 3 = 7 digit,
4th digit = 7 + 4 = 11 = 8 digit,
The next number should be 11.
BTW It is the lottery number of actor Innocent from the Malayalam movie Kilukkam

Answer by ikleyn(52788)

.
For other solutions look into these links

1)  https://www.quora.com/What-is-the-next-number-in-the-sequence-7-2-0-7-8-4-3

     https://www.quora.com/What-is-the-next-number-in-the-sequence-7-2-0-7-8-4-3



2)  https://www.algebra.com/algebra/homework/Sequences-and-series/Sequences-and-series.faq.question.1070387.html

     https://www.algebra.com/algebra/homework/Sequences-and-series/Sequences-and-series.faq.question.1070387.html